# Opening Articulated Objects in the Real World

### v2

@@ Abstract
What does it take to build mobile manipulation systems that can competently operate on previously unseen objects in previously unseen environments? This work answers this question using opening of articulated objects as a mobile manipulation testbed. Specifically, our focus is on the end-to-end performance on this task without any privileged information, i.e. the robot starts at a location with the novel target articulated object in view, and has to approach the object and successfully open it. We first develop a system for this task, and then conduct 100+ end-to-end system tests across 13 real world test sites. Our large-scale study reveals a number of surprising findings: a) modular systems outperform end-to-end learned systems for this task, even when the end-to-end learned systems are trained on 1000+ demonstrations, b) perception, and not precise end-effector control, is the primary bottleneck to task success, and c) state-of-the-art articulation parameter estimation models developed in isolation struggle when faced with robot-centric viewpoints. Overall, our findings highlight the limitations of developing components of the pipeline in isolation and underscore the need for system-level research, providing a pragmatic roadmap for building generalizable mobile manipulation systems. Videos, code, and models are available on the project website: https://arjung128.github.io/opening-articulated-objects/.
@@ I Introduction
To tackle the challenge of strong generalization within the context of mobile manipulation, we work on the task of opening articulated objects in diverse previously unseen real world environments (Figure LABEL:fig:dummy-teaser). Articulated objects, such as cabinets, drawers, and ovens, are ubiquitous in indoor environments, making their reliable manipulation a key capability for robots operating in the wild. In our problem setting, a Stretch RE2 robot (a commodity mobile manipulator without any hardware modifications) is placed in a novel environment in front of a previously unseen articulated object, and the objective is to solve the full end-to-end task. This includes detecting the object and estimating its articulation parameters, generating a precise whole-body motion plan, navigating to an optimal interaction position, securing a firm grasp on the handle, and successfully opening the object without collisions—all performed zero-shot, without any privileged information about the object or environment. Thus, this problem setting includes all of the key challenges that a mobile manipulator would face when interacting with previously unseen objects in unfamiliar environments in the real world, and thus forms a good test bed for study. We first developed an end-to-end mobile manipulation system to tackle this task and then conducted extensive end-to-end testing to understand the current bottlenecks in building such a system. This paper describes the system we build, the testing we conducted, and the lessons we learned along the way.
The System. We considered two broads ways of putting together such a system: a modular approach and an end-to-end learning approach, but ultimately favored a modular approach. Our approach, called MOSART for a MOdular System for opening ARTiculated objects, develops and utilizes state-of-the-art modules for perception, planning, and adaptation (shown in Figure I). Very briefly, MOSART adapts a Mask RCNN model [23] for inferring articulation parameters, extends a trajectory optimization framework for producing whole body motion plans [21], and utilizes proprioceptive feedback for mitigating last centimeter errors.
@@
• A modular system outperforms an end-to-end learning system when tested for generalization to unseen objects in novel environments. We compare MOSART to the recent and contemporary imitation learning work Robot Utility Models (RUM) from Etukuru et al. [12]. RUM is trained on a large-scale dataset of expert demonstrations: 1200 demos for opening cabinets, and 525 demonstrations for opening drawers across approximately 40 environments, by far the largest imitation learning dataset for this task. To our surprise, we find that the modular system outperforms this latest end-to-end learning method (Section IV-B and Table I). This results is particularly useful in context of the ongoing debate between modular learning and end-to-end learning in robotics at large, and specifically for recent end-to-end learning works that tackle mobile manipulation problems [53, 63, 65, 20].
• State-of-the-art perception modules for articulation parameter prediction struggle on images from a robot. Significant progress has been made in predicting articulation parameters from images [57, 29, 7, 48, 40, 61, 45, 41]. However, most of this progress has been made in isolation, without evaluating these methods as part of a complete, end-to-end system on real-world, in-the-wild images. When tested on viewpoints from the robot in the real world, recent state-of-the-art approaches such as OPD-Multi [57] and AO-Grasp [41] struggle to make accurate predictions. In contrast, our choice to develop a Articulation-parameter Prediction Module (APM) targeted towards images that the robot was expected to see, achieves substantially better performance vs. OPD-Multi [57] (Section IV-C1 and Figure 4) and AO-Grasp [41] (Section IV-C1, Table III, Figure A16).
@@
Broader Implications. Our findings provide key insights for designing generalizable mobile manipulation systems, offering several takeaways that should inform future research and system development. In particular, our large-scale, real-world experiments show that vision models trained in isolation frequently underperform when confronted with the distinct viewpoints encountered by robots, highlighting the need for comprehensive, system-level research. Moreover, we identify perception as the prsincipal bottleneck in achieving consistent task success, a finding that underscores the need for future research to prioritize robust perception modules. Additionally, our results indicate that modular approaches tend to outperform end-to-end methods, cautioning practitioners that scaling imitation-learning datasets to even 1,000 demonstrations does not necessarily lead to broad generalization. Notably, modular systems also facilitate easier adaptation to related tasks without requiring extensive new data collection. Finally, last-centimeter grasping remains a key challenge. While contact correction helps alleviate this failure mode to some extent, developing closed-loop interaction strategies could improve performance in tasks requiring precision, making it an important direction for future research. Overall, these lessons provide a pragmatic roadmap for researchers and practitioners aiming to build generalizable mobile manipulation systems.
@@ II-A Predicting Articulation Parameters
Researchers have looked at different aspects: a) construction of datasets (from simulation [33, 10, 16], real world images [57, 29, 7], and real world 3D scans [21, 58]), b) use of different input modalities to make predictions (RGB images [57], point clouds [36, 44, 60, 11, 69, 35, 51], RGB-D images [29, 64, 1, 68]), c) use of diverse sources of supervision [48], and d) predicting sites for interaction (i.e. handles) in addition to articulation parameters [40, 61, 45, 67]. Most past works are largely evaluated in simulation, and because no prediction is made for other articulation parameters (e.g. radius), these works cannot be used to generate motion plans to fully articulate objects. Sun et al. [57] add additional heads to Mask RCNN to predict articulation parameters in 3D. We also add additional heads to Mask RCNN but rather than directly predicting 3D outputs using RGB-D input, we take a 2D prediction from RGB and 3D lifting using depth image approach [22, 62].
@@ II-B Generating Motion Plans
Opening doors and drawers requires the end-effector to conform to the constraint defined by the object handle’s trajectory (which in turn is determined by the location of the articulation joint and the handle). Depending on the robot morphology, robot configurations that conform to this end-effector constraint might represent a measure zero set. This makes it challenging to directly use sampling-based motion planners [31, 34]. Past research has therefore developed specialized methods for planning under such constraints [32, 4] and used it to articulate objects [5, 9, 50, 47, 38, 6, 59, 24, 25, 42]. Another line of work casts it as trajectory optimization [21, 71, 52] or optimal control  [13, 46, 54, 39] problem. As these only search for a solution locally (vs. motion planning that searches globally via sampling), it is important to properly initialize the trajectory optimizers. Recent work from Gupta et al. [21] alleviates this limitation by using learning to predict good initializations for trajectory optimization thus generating high-quality motion plans quickly. We adopt their approach but extend it to produce whole body motion plans as we describe in Section III.
@@ III Modular System for Opening Articulated Objects (MOSART)
We follow a modular approach comprising of a) a Perception Module that predicts articulation parameters using on-board an RGB-D camera, b) a Motion Plan Generator that converts predicted articulations into whole-body motion plans, and c) an Execution Module that adapts and executes the generated motion plan using proprioceptive feedback. Figure I shows an overview. We describe the pipeline in terms of two prototypical articulated objects: drawers for prismatic joints and cabinets for hinged joints, but our pipeline is more general as our experiments will reveal.
@@ III-C Adapting and executing motion plans using proprioceptive feedback
Minor errors in state estimation, navigation, and calibration compound to prevent handle grasping. Specifically, there is little tolerance in the depth direction. This is further exacerbated by the hardware design of the gripper that becomes shorter upon closure. We combat this via a contact-based correction strategy for arriving at the pre-grasp pose: we extend the gripper towards the object until contact is made. See Figure 2 for a visualization of the strategy for the different articulated objects. As is, the end-effector approaches the handles from the side, which works well for horizontal handles, but poses issues for vertical handles. For objects with vertical handles, we thus have an additional vertical primitive before the contact-based correction mechanism. We use the arm effort signal and the gripper’s yaw effort signal to detect contact. The change in the end-effector pose during contact correction phase provides feedback to update the desired trajectory and motion plan. We ablate the use of contact correction in Section IV-C2.
@@ III-D Full end-to-end execution
Starting from where the robot can see the target articulated object, we use APM to predict the handle location (in the camera’s coordinate frame), the surface normal, and the radius (for cabinets). We use a calibrated robot URDF to transform these 3D predictions from the camera frame to the base frame. Next, we generate a whole-body motion plan and execute the first qpose (full robot configuration), followed by our contact-based correction mechanism. Once the handle has been grasped, we execute the rest of the motion plan.
IV Experiments
We work with the Stretch RE2 robot. We first present our end-to-end system test results, evaluating MOSART on 31 novel drawers and cupboards across 10 buildings (Section IV-A). To see how a modular system compares to an end-to-end learning approach, we compare MOSART to RUM [12] and a sim2real imitation learning approach (Section IV-B). To further understand the system, we then evaluate individual pipeline modules (Section IV-C). This includes evaluating the quality of our MaskRCNN-based perception module (as well as a Detic-based perception model) on real world images, comparing APM to two recent articulation parameter prediction systems [41, 57], and an ablation for the contact-based correction mechanism. To understand the interaction between perception and control, we quantify the robustness of execution to state estimation inaccuracies (Section IV-D). We then study the generalization of our pipeline to other articulation types and diverse handles (Section IV-E), before we analyzing the failure modes of our system (Section IV-F).
IV-A End-to-end System Tests
@@ IV-B1 Robot Utility Models (RUM)
As shown in Table I, MOSART substantially outperforms RUM in a paired evaluation on unseen objects across four novel environments. One common failure mode we observe for RUM is the inability to grasp the handle of the object, despite getting very close to it. Another is incorrect handle selection, particularly when multiple handles are in close proximity and within view. This stems from RUM’s inability to explicitly specify the task, making it susceptible to such ambiguities and highlighting a key limitation of the approach. Furthermore, our evaluation focuses solely on in-the-wild objects, where such challenging cases naturally arise, whereas a considerable portion of RUM’s evaluation was conducted on isolated objects in controlled lab environments. These findings highlight a) the importance of end-to-end system tests in-the-wild, b) the challenging nature of this problem that requires generalization to previously unseen objects in novel environments, and c) the effectiveness of MOSART.
@@ IV-C1 Articulation Parameter Prediction Accuracy
We also compare APM to two recent works: AO-Grasp [41] and OPDMulti [57]. We perform a quantitative comparison to AO-Grasp. Table III reports a quantitative comparison of the mean handle error on our real world images (lower is better). Our method significantly outperforms AO-Grasp. See Section A12 for a qualitative comparison to AO-Grasp, and for more details. We also perform a qualitative comparison to OPDMulti (see Figure 4), and find that our model produces more accurate segmentation masks (which affect the surface normal, and thus the navigation). OPDMulti does not predict handle locations, disqualifying it as a replacement to APM, and preventing us from performing a quantitative comparison. Additionally, we also compare to an approach which directly predicts articulation parameters in 3D (Section A13).
@@
Table IV: Ablations. Success rates for our full pipeline with and without contact correction (across all objects) in the real world. Contact correction significantly improves the success rate.
@@ V Limitations
While we evaluated MOSART on a variety of objects, our study excluded those with round knobs or no handles. MOSART uses pre-mined navigation targets without allowing for base translation, which may limit performance on objects surrounded by challenging collision geometry or with large-radii. Lastly, the mapping from articulation parameters to waypoints must be provided to MOSART, whereas an end-to-end learned approach may learn this from data.

### v3

# Opening Articulated Structures in the Real World

@@ Abstract
What does it take to build mobile manipulation systems that can competently operate on previously unseen objects in previously unseen environments? This work answers this question using opening of articulated structures as a mobile manipulation testbed. Specifically, our focus is on the end-to-end performance on this task without any privileged information, i.e. the robot starts at a location with the novel target articulated object in view, and has to approach the object and successfully open it. We first develop a system for this task, and then conduct 100+ end-to-end system tests across 13 real world test sites. Our large-scale study reveals a number of surprising findings: a) modular systems outperform end-to-end learned systems for this task, even when the end-to-end learned systems are trained on 1000+ demonstrations, b) perception, and not precise end-effector control, is the primary bottleneck to task success, and c) state-of-the-art articulation parameter estimation models developed in isolation struggle when faced with robot-centric viewpoints. Overall, our findings highlight the limitations of developing components of the pipeline in isolation and underscore the need for system-level research, providing a pragmatic roadmap for building generalizable mobile manipulation systems. Videos, code, and models are available on the project website: https://arjung128.github.io/opening-articulated-structures/.
@@ I Introduction
To tackle the challenge of strong generalization within the context of mobile manipulation, we work on the task of opening articulated structures in diverse previously unseen real world environments (Figure LABEL:fig:dummy-teaser). Articulated structures, such as cabinets, drawers, and ovens, are ubiquitous in indoor environments, making their reliable manipulation a key capability for robots operating in the wild. In our problem setting, a Stretch RE2 robot (a commodity mobile manipulator without any hardware modifications) is placed in a novel environment in front of a previously unseen articulated object, and the objective is to solve the full end-to-end task. This includes detecting the object and estimating its articulation parameters, generating a precise whole-body motion plan, navigating to an optimal interaction position, securing a firm grasp on the handle, and successfully opening the object without collisions—all performed zero-shot, without any privileged information about the object or environment. Thus, this problem setting includes all of the key challenges that a mobile manipulator would face when interacting with previously unseen objects in unfamiliar environments in the real world, and thus forms a good test bed for study. We first developed an end-to-end mobile manipulation system to tackle this task and then conducted extensive end-to-end testing to understand the current bottlenecks in building such a system. This paper describes the system we build, the testing we conducted, and the lessons we learned along the way.
The System. We considered two broad ways of putting together such a system: a modular approach and an end-to-end learning approach, but ultimately favored a modular approach. Our approach, called MOSART for a MOdular System for opening ARTiculated structures, develops and utilizes state-of-the-art modules for perception, planning, and adaptation (shown in Figure I). Very briefly, MOSART adapts a Mask RCNN model [28] for inferring articulation parameters, extends a trajectory optimization framework for producing whole body motion plans [26], and utilizes proprioceptive feedback for mitigating last centimeter errors.
@@
• A modular system outperforms an end-to-end learning system when tested for generalization to unseen objects in novel environments. We compare MOSART to the recent and contemporary imitation learning work Robot Utility Models (RUM) from Etukuru et al. [16]. RUM is trained on a large-scale dataset of expert demonstrations: 1200 demos for opening cabinets, and 525 demonstrations for opening drawers across approximately 40 environments, by far the largest imitation learning dataset for this task. To our surprise, we find that the modular system outperforms this latest end-to-end learning method (Section IV-B and Table I). This result is particularly useful in the context of the ongoing debate between modular learning and end-to-end learning in robotics at large, and specifically for recent end-to-end learning works that tackle mobile manipulation problems [72, 82, 84, 25].
• State-of-the-art perception modules for articulation parameter prediction struggle on images from a robot. Significant progress has been made in predicting articulation parameters from images [76, 36, 11, 63, 52, 80, 60, 53]. However, most of this progress has been made in isolation, without evaluating these methods as part of a complete, end-to-end system on real-world, in-the-wild images. When tested on viewpoints from the robot in the real world, recent state-of-the-art approaches such as OPD-Multi [76] and AO-Grasp [53] struggle to make accurate predictions. In contrast, our choice to develop an Articulation-parameter Prediction Module (APM) targeted towards images that the robot was expected to see, achieves substantially better performance vs. OPD-Multi [76] (Section IV-C1 and Figure 4) and AO-Grasp [53] (Section IV-C1, Table III, Figure A16).
@@
Broader Implications. Our findings provide key insights for designing generalizable mobile manipulation systems, offering several takeaways that should inform future research and system development. In particular, our large-scale, real-world experiments show that vision models trained in isolation frequently underperform when confronted with the distinct viewpoints encountered by robots, highlighting the need for comprehensive, system-level research. Moreover, we identify perception as the principal bottleneck in achieving consistent task success, a finding that underscores the need for future research to prioritize robust perception modules. Additionally, our results indicate that modular approaches tend to outperform end-to-end methods, cautioning practitioners that scaling imitation-learning datasets to even 1,000 demonstrations does not necessarily lead to broad generalization. Notably, modular systems also facilitate easier adaptation to related tasks without requiring extensive new data collection. Finally, last-centimeter grasping remains a key challenge. While contact correction helps alleviate this failure mode to some extent, developing closed-loop interaction strategies could improve performance in tasks requiring precision, making it an important direction for future research. Overall, these lessons provide a pragmatic roadmap for researchers and practitioners aiming to build generalizable mobile manipulation systems.
@@ II-A Predicting Articulation Parameters
Researchers have extensively looked at different aspects: a) construction of various datasets (from simulation [40, 14, 20], real world images [76, 36, 11], and real world 3D scans [26, 77]), b) use of different input modalities to make predictions (RGB images [76], point clouds [44, 59, 79, 15, 88, 42, 70], RGB-D images [36, 83, 1, 87]), c) use of diverse sources of supervision [63], and d) predicting sites for interaction (i.e. handles) in addition to articulation parameters [52, 80, 60, 86]. Most past works are largely evaluated in simulation, and because no prediction is made for other articulation parameters (e.g. radius), these works cannot directly be used to generate motion plans to fully articulate objects in the real world without any privileged information. Sun et al. [76] augment the Mask RCNN architecture by adding additional heads specifically designed to predict articulation parameters in 3D. We also add additional heads to Mask RCNN; however, rather than directly predicting 3D outputs from the RGB-D input, we adopt a two-stage approach involving 2D prediction from RGB images followed by 3D lifting using the depth image [27, 81].
@@ II-B Generating Motion Plans
Opening doors and drawers requires the end-effector to conform to the constraint defined by the object handle’s trajectory (which in turn is determined by the location of the articulation joint and the handle). Depending on the robot morphology, robot configurations that conform to this end-effector constraint might represent a measure zero set. This makes it challenging to directly use sampling-based motion planners [38, 41]. Past research has therefore developed specialized methods for planning under such constraints [39, 4] and used it to articulate objects [5, 13, 68, 62, 49, 6, 78, 30, 31, 57]. Another line of work casts it as a trajectory optimization [26, 90, 71] or optimal control  [17, 61, 73, 51] problem. As these only search for a solution locally (vs. motion planning that searches globally via sampling), it is important to properly initialize the trajectory optimizers. Recent work from Gupta et al. [26] alleviates this limitation by using learning to predict good initializations for trajectory optimization thus generating high-quality motion plans quickly. We adopt their approach but extend it to produce whole body motion plans as we describe in Section III.
@@ II-C Mobile Manipulation
II-D Modular Learning
Methods that integrate learning-based components with classical methods, have emerged as an effective approach for building robust robotic systems. Examples span a wide variety of domains: grasping [55, 47, 54, 46], aerial robotics [69], and autonomous driving [56, 48]. For instance, DexNet 2.0 [47] uses a neural network to evaluate grasp quality and a sampling-based planner to execute the chosen grasp. A large body of work in robotic navigation also utilizes modular learning [9, 8, 64, 10, 65, 7, 22]. Notably, sharing similarities with our work, Gervet et al. [22] design a modular system for navigation, and via a large real world evaluation, demonstrate the superior performance of such a system relative to classical approaches and end-to-end learning. Recently, modular learning approaches have also leveraged vision-language models (VLMs) [32] and have been applied successfully to humanoid control [29, 43]. For example, ReKep [32] employs vision foundation models for high-level reasoning, followed by trajectory optimization for low-level control. While modular learning methods vary in the choice of modules and their reliance on learning, in this work, we present a practical and effective modular design demonstrated on the challenging mobile manipulation task of opening articulated objects in-the-wild.
@@ III Modular System for Opening Articulated Objects (MOSART)
We follow a modular approach comprising of a) a Perception Module that predicts articulation parameters using an on-board RGB-D camera, b) a Motion Plan Generator that converts predicted articulations into whole-body motion plans, and c) an Execution Module that adapts and executes the generated motion plan using proprioceptive feedback. Figure I shows an overview. We describe the pipeline in terms of two prototypical articulated objects: drawers for prismatic joints and cabinets for hinged joints, but our pipeline is more general as our experiments will reveal.
@@ III-C Adapting and executing motion plans using proprioceptive feedback
Minor errors in state estimation, inaccuracies in navigation, and imperfections in calibration can compound over time to ultimately prevent successful handle grasping. Specifically, the system exhibits very little tolerance to errors along the depth direction. This is further exacerbated by the hardware design of the gripper that becomes shorter upon closure. To combat this compounded error, we employ a contact-based correction strategy aimed at refining the pre-grasp pose: specifically, we extend the gripper incrementally towards the object surface until physical contact is detected. See Figure 2 for a visualization of the strategy for the different articulated objects. As is, the end-effector approaches the handles from the side, which works well for horizontal handles, but poses issues for vertical handles. For objects with vertical handles, we thus have an additional vertical primitive before the contact-based correction mechanism. We use the arm effort signal and the gripper’s yaw effort signal to detect contact. The change in the end-effector pose during contact correction phase provides feedback to update the desired trajectory and motion plan. We ablate the use of contact correction in Section IV-C2.
@@ III-D Full end-to-end execution
Starting from where the robot can see the target articulated object, we use APM to predict several key properties: the handle location (in the camera’s coordinate frame), the surface normal of the surface, and the articulation radius (for cabinets). We use a calibrated robot URDF to transform these 3D predictions from the camera frame to the base frame. Next, we generate a whole-body motion plan and execute the first qpose (full robot configuration), followed by our contact-based correction mechanism. Once the handle has been grasped, the robot proceeds to execute the remainder of the precomputed motion plan to complete the task.
@@
IV Experiments
We work with the Stretch RE2 robot. We first present our end-to-end system test results, evaluating MOSART on 31 novel drawers and cupboards across 10 buildings (Section IV-A). To see how a modular system compares to an end-to-end learning approach, we compare MOSART to RUM [16] and a sim2real imitation learning approach (Section IV-B). To further understand the system, we then evaluate individual pipeline modules (Section IV-C). This includes evaluating the quality of our MaskRCNN-based perception module (as well as a Detic-based perception model) on real world images, comparing APM to two recent articulation parameter prediction systems [53, 76], and an ablation for the contact-based correction mechanism. To understand the interaction between perception and control, we quantify the robustness of execution to state estimation inaccuracies (Section IV-D). We then study the generalization of our pipeline to other articulation types and diverse handles (Section IV-E), before we analyze the failure modes of our system (Section IV-F).
IV-A End-to-end System Tests
@@ IV-B1 Robot Utility Models (RUM)
As shown in Table I, MOSART substantially outperforms RUM in a paired evaluation on unseen objects across four novel environments. In addition to achieving higher success rates, MOSART is also faster than RUM (see Section A13). One common failure mode we observe for RUM is the inability to grasp the handle of the object, despite getting very close to it. Another is incorrect handle selection, particularly when multiple handles are in close proximity and within view. This stems from RUM’s inability to explicitly specify the task, making it susceptible to such ambiguities and highlighting a key limitation of the approach. Furthermore, our evaluation focuses solely on in-the-wild objects, where such challenging cases naturally arise, whereas a considerable portion of RUM’s evaluation was conducted on isolated objects in controlled lab environments. These findings highlight a) the importance of end-to-end system tests in-the-wild, b) the challenging nature of this problem that requires generalization to previously unseen objects in novel environments, and c) the effectiveness of MOSART.
@@ IV-C1 Articulation Parameter Prediction Accuracy
In addition to our primary evaluations, we also compare APM to two recent works: AO-Grasp [53] and OPDMulti [76]. We perform a direct quantitative comparison to AO-Grasp to assess differences in handle prediction accuracy. Table III reports a quantitative comparison of the mean handle error on our real world images (lower is better). Our method significantly outperforms AO-Grasp. See Section A14 for a qualitative comparison to AO-Grasp, and for more details. We also perform a qualitative comparison to OPDMulti (see Figure 4), and find that our model produces more accurate segmentation masks (which affect the surface normal, and thus the navigation). OPDMulti does not predict handle locations, disqualifying it as a replacement to APM, and preventing us from performing a quantitative comparison. Additionally, we also compare to an approach which directly predicts articulation parameters in 3D (Section A15), as well as to Grounded SAM 2 (Section A16).
@@
Table IV: Ablations. Success rates for our full pipeline, both with and without contact correction, evaluated across all objects in the real world. Incorporating contact correction significantly improves the success rate.
@@ V Limitations
While we evaluated MOSART on a variety of objects, our study excluded those with round knobs or no handles. MOSART uses pre-mined navigation targets without allowing for base translation, which may limit performance on objects surrounded by challenging collision geometry or with large-radii. Additionally, the mapping from articulation parameters to waypoints must be provided to MOSART, whereas an end-to-end learned approach may learn this from data. Finally, there are limitations of the embodiment we use (e.g. it cannot reach cabinets high up, or exert enough force to pull open fridge doors).